OpenStudy (anonymous):

Determine the constant of variation -x^2/y=3

OpenStudy (anonymous):

@jdoe0001

jimthompson5910 (jim_thompson5910):

Solve for y to get -x^2/y=3 -x^2 = 3y 3y = -x^2 y = (-x^2)/3 y = (-1/3)x^2

jimthompson5910 (jim_thompson5910):

The equation is in the form y = k*x^2 where in this case, the constant of variation is k = -1/3